# Efficient Modelling of Lyman-$\alpha$ opacity fluctuations during late EoR

### v1

@@ 4 Observational Data & Likelihood Analysis
We utilize the high quality Ly-$\alpha$ forest data from 67 high redshift quasars following Bosman et al. (2022). This sample includes 51 optical quasar spectra from VLT X-shooter spectrograph, among which 25 are from XQR-30 program (D’Odorico et al., 2023) and 26 are from archival X-shooter data. The remaining 16 spectra have been taken with the Keck ESI spectrograph. All these high redshift (at a redshift $z\geq 5.6$) quasars spectra with SNR $\geq 10$ per pixel provide an ideal dataset to probe the late phase of the reionization epoch. However, the quasar continuum is very uncertain at these redshifts due to the lack of enough pixels to recover the continuum flux. Hence, to compute effective optical depth, a novel principal component analysis (PCA) method has been followed in recent studies (Davies et al., 2018c; Bosman et al., 2021). Then the effective optical depths are computed with the bins of $\Delta z=0.1$ (Bosman et al., 2022).
In our analysis, the observational uncertainties and wavelength dependent continuum uncertainties are taken into account via forward modelling. To be specific, the measurement uncertainties are added with each realization of the model skewers according to the data, while the continuum uncertainties are multiplicative. The sightlines for which no significantly detected transmission is available, are treated as optimistic limits i.e. assumed to be $2\sigma$ where $\sigma$ is the data uncertainty corresponding to that sightline. In pessimistic scenario, the flux could be assumed to be zero producing infinitely opaque sightline.
Given the above dataset, we use the standard Bayesian approach for parameter space exploration. Specifically, we compute the conditional probability distribution or the posterior $\mathcal{P}(\theta|\mathcal{D})$ of the model parameters $\theta$, provided the observational data sets $\mathcal{D}$ as mentioned in the previous paragraph. This can be estimated using the Bayes theorem
$\mathcal{P}(\theta|\mathcal{D})=\frac{\mathcal{L}(\mathcal{D}|\theta)\leavevmode\nobreak\ \pi(\theta)}{\mathcal{P}(\mathcal{D})},$ (9)
@@
However, the raw dataset can not be used for likelihood analysis due to the dominance of Poisson noise arising from its discrete nature. To this end, different summary statistics, such as Cumulative Distribution Function (CDF), mean fluctuations, etc can be used for inference studies with the assumption that the posterior $\mathcal{P}(\theta|\mathcal{D})\approx\mathcal{P}(\theta|\mathcal{D_{S}})$, where $\mathcal{D_{S}}$ is the corresponding statistics derived from data. Among these, Probability Distribution Functions (PDFs) fitted with Gaussian Kernel Density Estimator (KDE) are commonly used Gaikwad et al. (2021); Choudhury et al. (2021) for robust likelihood inference. Here, we adopt similar approach and evaluate the KDE distribution into 15 bins to compute the likelihood. We do the exact same procedure with both the observational data and our models. We keep the number of sightlines same as the data. Once we have the binned probability distribution, we can compute the logarithm of likelihood ($\mathcal{L}$) distribution as
@@
where $p_{\mathcal{D}}(\tau_{\mathrm{eff}})$ is the observed KDE PDF of effective optical depth and $p_{\mathcal{M}}(\tau_{\mathrm{eff}})$ is the corresponding model KDE PDF. We also take into account for the correlations between the bins by computing the covariance matrix $\Sigma_{ij}(\mathcal{M})$, which is model dependent. This is computed using 1000 random independent realizations of the sightlines.
At the left panel of Figure 9, we show the probability distributions of the data along with the fiducial model at redshift, $z=5.4$. The blue histogram represents the effective optical depth distribution according to the original data while the black dashed line is achieved after applying the Gaussian KDE on it. The estimation from the model is shown in red which seems to match the data KDE reasonably well. At the right panel, we show the normalized covariance matrix for this model. Notably, our covariance matrices are dependent upon the model which further goes into the computation of likelihood for MCMC sampling.
Next, we sample the posterior distribution using the Monte Carlo Markov Chain (MCMC) method, more specifically, the Metropolis-Hastings algorithm (Metropolis et al., 1953). We utilize the publicly available package cobaya (Torrado & Lewis, 2021) [^1] https://cobaya.readthedocs.io/en/latest/ to run the MCMC chains. We check the convergence of the chains following the Gelman-Rubin $R-1$ statistic (Gelman & Rubin, 1992). The chain is assumed to be converged if the $R-1$ value becomes less than a threshold $0.01$.
Figure 13: A comparison of UVB fluctuation fields allowed by the posteriors at redshift, $z=5.8$ (reionization is end) and $z=5.9$ (reionization is incomplete). The mean free path requirement for $z=5.8$ ($\lambda_{0}=4\leavevmode\nobreak\ h^{-1}\mathrm{cMpc}$) is much smaller than $z=5.9$ ($\lambda_{0}=40\leavevmode\nobreak\ h^{-1}\mathrm{cMpc}$), showing the effect of incomplete ionization on the UVB field.
Figure 14: Marginalized 1D posterior constraints for and mean photoionization rate ($\langle\Gamma_{12}^{\mathrm{HI}}\rangle$) and effective mean free path ($\lambda_{0}$) at different redshifts. The constraints correspond to 95% confidence limits where the constraints are available from both sides. For limits, we quote the value where posterior falls down to 1/$e^{2}$ from the maximum. The red circles denote the constraints derived from this work (hollow circles represent the redshift where reionization is not complete). The redshifts for green and blue points have been shifted by 0.02 dex for visual clarity.
@@ 5.1 Cumulative probability distribution of $\tau_{\mathrm{eff}}$
We show the cumulative probability distributions (CDFs) of the effective optical depth ($\tau_{\mathrm{eff}}$) in Figure 10, for the redshift range of our interest ($z=5.0-6.1$) in this study. The red lines show the bestfit CDF obtained by combining 500 random realizations of the sightlines while the number of sightlines is kept the same as provided by the observational data. The cyan curves in the background represent the individual realizations of the bestfit model. For the data (shown in black dashed lines), we show both optimistic and pessimistic limits although we use only the optimistic case for the parameter space exploration. We check that the bestfit evolution model can nicely match the observed scatter in the optical depth distribution as well as the mean transmission flux in majority of the cases with mild deviation at $z=5.6$ and $5.9$. The deviation is mainly due to the dominance of $\tau_{\mathrm{eff}}$ values at the distribution of the edges in the data for those redshifts. Furthermore, the width of the distribution increases as we go towards higher redshifts suggesting increasing fluctuations in the opacity fields.
@@ 5.2 Joint parameter posteriors
Next, we quantify the posterior distribution of the parameters extracted from three different temperature models allowed by the current constraints on the IGM temperatures as shown earlier in Figure 5. We perform this analysis for all the redshifts of interests and show the combined posteriors in Figure 11 while in Figure 12, we show the 2D posterior contours (at 68% and 95% confidence level) for the individual temperature models along with the combined contour of all the three samples again at the fiducial redshift, $z=5.4$. It is apparent that the combined contour nicely represents the whole posterior space allowed by the temperature constraints. Now, moving the discussion with Figure 11 for different redshifts, it can be seen that the data are consistent even with almost uniform UVB models towards low redshifts ($z\leq 5.3$), demanding only a small amount of UVB fluctuations. An interesting feature can be noticed at $z=5.3$ where a bimodality in mean free path distribution arises disfavouring some of the moderate UVB fluctuation models ($\lambda_{0}\sim 20-30h^{-1}\mathrm{cMpc}$). This happens because those models are not able to provide enough fluctuations to match the data distribution. For a large mean free path, the opacity fluctuations are driven by the underlying density field and temperature fluctuations. As the mean free path decreases, it initially neutralizes the opacity fluctuations as demonstrated in Figure 7, which corresponds to the disfavored region in the posterior. This happens because high density regions are ionized earlier and have relatively lower temperatures at later times due to availability of more time for cooling (D’Aloisio et al., 2015). This enhances the opacity fluctuations correlating with density fluctuations. However, the correlation between density and UVB field is opposite providing high photoionization rate at high density regions and vice versa which further drives the fluctuations on the other direction (see also; Davies et al., 2018a). Then, with a reasonably short mean free path, the opacity fluctuations are dominated by the UVB field itself. Hence, the bimodal feature portrays the interplay among these competing fluctuations generating the opacity field.
@@
As we move towards higher redshifts (up to $z=5.8$), where reionization ends in our model), the required UVB fluctuations also increase allowing for shorter mean free paths. Before reionization ends ($z\geq 5.9$), the ionization fluctuation is strong enough to produce the required opacity fluctuations, hence requiring little or almost no fluctuations from the UVB field. This further keeps the mean free path parameter unconstrained. This explanation can be more clear from Figure 13, where we show two snapshots of UVB fluctuations. Among these two snapshots, one is at redshift, $z=5.8$ (after reionization completion) with a very short mean free path ($\lambda_{0}=4\leavevmode\nobreak\ h^{-1}\mathrm{cMpc}$) and the other one is at redshift, $z=5.9$ (before reionization completion) with a relatively high mean free path parameter ($\lambda_{0}=40\leavevmode\nobreak\ h^{-1}\mathrm{cMpc}$), as suggested by the posteriors. It is apparent that the presence of neutral regions can produce strong fluctuations in the UVB field even if $\lambda_{0}$ is larger (see also, Choudhury et al., 2021). This supports our above argument for an unconstrained mean free path at higher redshifts where reionization is still not ended. This analysis provides us the confidence that our simplistic model can be utilized to pursue parameter space statistics study and interpret the relevant physics from the observational data.
@@
In Figure 14, we show the marginalized constraints on the mean photoionization rate ($\langle\Gamma_{12}^{\mathrm{HI}}\rangle$, at the left panel) and effective mean free path ($\lambda_{0}$, at the right panel) at 95% confidence level along with the points from recent literature. Our estimates are denoted with red circles (the solid ones indicate the redshifts where reionization is completed and hollow ones indicate the redshifts with incomplete reionization). We found that our photoionization rate estimates are consistent with the findings of Davies et al. (2024b, shown in blue stars with displaced redshift for better visualization). This is not very surprising as the UVB fluctuation generation model is the same for both cases. However, our points are slightly higher as we have a realistic reionization history allowing sufficient time for cooling in the early ionized high density regions. On the other hand, Davies et al. (2024b) assumed a sudden reionization for all the cells in the simulation box which results in larger temperature and hence, a slightly shorter photoionization rate to match the mean transmission flux. Additionally, our estimates are also in agreement with the corresponding constraints of Gaikwad et al. (2023, shown in green triangle again shifted in redshift).
@@
In case of the mean free path, we are able to derive the constraints from posteriors at redshift range $z=5.4-5.7$. For the redshifts providing constraints from only one side, we quote the value where the posterior value falls down to $1/e^{2}$ from the maximum (Davies et al., 2024b). We do not get any constraints at $z=6.0$. Our constraints prefer a shorter mean free path distribution than Davies et al. (2024b) values, although these are consistent within the errorbars, particularly given that the lower end of the Davies et al. (2024b) uncertainties correspond to an approximate correction to account for their lack of model self-consistency. The main reason for this difference is again traced back to the assumed sudden reionization and the lack of post reionization large scale temperature fluctuations in their model. It is expected that the large scale temperature field is anticorrelated with the density field due to the variations in reionization timing at different regions. As a result of that a stronger UVB fluctuation is required to overcome the correlated fluctuations of thermal and density field which in turn demands a lower $\lambda_{0}$ value. Intriguingly, this more physical treatment with a gradual and patchy thermal evolution puts our model constraints within the similar ballpark of recent high redshift mean free path measurements (Becker et al., 2021; Zhu et al., 2023). At this point, it is important to highlight that the definition of our mean free path is similar to Davies et al. (2024b), which is solely a subgrid parameter quantifying clumping of the unresolved gas. Hence, our inferred mean free path only characterizes the excess UVB fluctuations over the uniform case without including density field fluctuations. As a consequence, we do not get any upper bound on $\lambda_{0}$ for $z\leq 5.3$ as an almost uniform UVB field is consistent with fluctuations required by the data. This is unlike the approach adopted by Gaikwad et al. (2023) where the mean free path is computed along with their density field after including fluctuating UVB contribution. That further enables them to put constraints on the mean free path even without requiring explicit UVB fluctuations. At redshift $z\geq 5.9$, reionization is incomplete in our model which further allows us to provide only lower limits on the mean free path, as demonstrated earlier in the subsection, 5.2. The constraints used in the plots are tabulated in Table 1.
@@
In continuation with the above discussion, it is worthwhile to mention that most of the semi numerical models give a sharp transition to the reionization end while detailed numerical simulations (Kannan et al., 2022; Asthana et al., 2025) prefer relatively smooth transitions along with slight inflection in neutral fraction evolution. A gradual transition will translate to a smoother evolution in mean free path as hinted from our findings. A smooth transition towards reionization end can be achieved within a semi-numerical set up with more realistic modelling of photon sinks (e.g. Davies & Furlanetto 2022; Qin et al. 2024). This can provide a direction for potential improvement for our model in the future.
@@
Table 1: Parameter constraints obtained from the MCMC-based analysis after combining different temperature models. For each redshift, we show the mean value of combined samples along with 95% confidence limits on different parameters where constraints are achieved from both side. We quote the value where posterior value falls down to $1/e^{2}$ from the maximum (Davies et al., 2024b) for one sided limit.

### v2

@@ 4 Observational data & likelihood analysis
We utilized the high quality Ly-$\alpha$ forest data from 67 high-redshift quasars following Bosman et al. (2022). This sample includes 51 optical quasar spectra from VLT X-shooter spectrograph, among which 25 are from the XQR-30 program (D’Odorico et al., 2023) and 26 are from archival X-shooter data. The remaining 16 spectra have been taken with the Keck ESI spectrograph. All these high redshift (at a redshift $z\geq 5.6$) quasars spectra with SNR $\geq 10$ per pixel provide an ideal dataset to probe the late phase of the reionization epoch. However, the quasar continuum is very uncertain at these redshifts due to the lack of enough pixels to recover the continuum flux. Hence, to compute effective optical depth, a novel principal component analysis (PCA) method has been followed in recent studies (Davies et al., 2018c; Bosman et al., 2021). Then the effective optical depths are computed with the bins of $\Delta z=0.1$ (Bosman et al., 2022).
In our analysis, the observational uncertainties and wavelength-dependent continuum uncertainties are taken into account via forward modelling. To be specific, the measurement uncertainties are added with each realization of the model skewers according to the data, while the continuum uncertainties are multiplicative. The sightlines for which no significantly detected transmission is available, are treated as optimistic limits, i.e., assumed to be $2\sigma$ where $\sigma$ is the data uncertainty corresponding to that sightline. In a pessimistic scenario, the flux could be assumed to be zero, producing an infinitely opaque sightline (following, Bosman et al., 2022).
Given the above dataset, we used the standard Bayesian approach for parameter space exploration. Specifically, we computed the conditional probability distribution or the posterior $\mathcal{P}(\theta|\mathcal{D})$ of the model parameters $\theta$, provided the observational data sets $\mathcal{D}$ as mentioned in the previous paragraph. This can be estimated using the Bayes theorem
$\mathcal{P}(\theta|\mathcal{D})=\frac{\mathcal{L}(\mathcal{D}|\theta)\penalty 10000\ \pi(\theta)}{\mathcal{P}(\mathcal{D})},$ (9)
@@
However, the raw dataset can not be used for likelihood analysis due to the dominance of Poisson noise arising from its discrete nature. To this end, different summary statistics, such as Cumulative Distribution Function (CDF), mean fluctuations, etc can be used for inference studies with the assumption that the posterior $\mathcal{P}(\theta|\mathcal{D})\approx\mathcal{P}(\theta|\mathcal{D_{S}})$, where $\mathcal{D_{S}}$ is the corresponding statistics derived from data. Among these, Probability Distribution Functions (PDFs) fitted with Gaussian Kernel Density Estimator (KDE) are commonly used Gaikwad et al. (2021); Choudhury et al. (2021) for robust likelihood inference. Here, we adopted a similar approach and evaluated the KDE distribution into 15 bins to compute the likelihood. We did the exact same procedure with both the observational data and our models. We kept the number of sightlines same as the data. Once we had the binned probability distribution, we computed the logarithm of the likelihood ($\mathcal{L}$) distribution as
@@
where $p_{\mathcal{D}}(\tau_{\mathrm{eff}})$ is the observed KDE PDF of effective optical depth and $p_{\mathcal{M}}(\tau_{\mathrm{eff}})$ is the corresponding model KDE PDF. We also took into account the correlations between the bins by computing the covariance matrix $\Sigma_{ij}(\mathcal{M})$, which is model dependent. This was computed using 1000 random independent realizations of the sightlines.
In the left panel of Figure 9, we show the probability distributions of the data along with the fiducial model at redshift, $z=5.4$. The blue histogram represents the effective optical depth distribution according to the original data, while the black dashed line is achieved after applying the Gaussian KDE on it. The estimation from the model is shown in red, which seems to match the data KDE reasonably well. In the right panel, we show the normalized covariance matrix for this model. Notably, our covariance matrices are dependent upon the model, which further goes into the computation of the likelihood for MCMC sampling.
Next, we sampled the posterior distribution using the Monte Carlo Markov Chain (MCMC) method, more specifically, the Metropolis-Hastings algorithm (Metropolis et al., 1953). We utilized the publicly available package cobaya (Torrado & Lewis, 2021) [^1] https://cobaya.readthedocs.io/en/latest/ to run the MCMC chains. We checked the convergence of the chains following the Gelman-Rubin R1R-1 statistic (Gelman & Rubin, 1992). The chain was assumed to be converged if the R1R-1 value became less than a threshold $0.01$.
Figure 13: A comparison of UVB fluctuation fields allowed by the posteriors at redshift, $z=5.8$ (reionization is end) and $z=5.9$ (reionization is incomplete). The mean free path requirement for $z=5.8$ ($\lambda_{0}=4\penalty 10000\ h^{-1}\mathrm{cMpc}$) is much smaller than $z=5.9$ ($\lambda_{0}=40\penalty 10000\ h^{-1}\mathrm{cMpc}$), showing the effect of incomplete ionization on the UVB field.
[…]
@@ 5.1 Cumulative probability distribution of $\tau_{\mathrm{eff}}$
We show the cumulative probability distributions (CDFs) of the effective optical depth ($\tau_{\mathrm{eff}}$) in Figure 10, for the redshift range of our interest (z=5.06.1z=5.0-6.1) in this study. The red lines show the bestfit CDF obtained by combining 500 random realizations of the sightlines while the number of sightlines is kept the same as provided by the observational data. The cyan curves in the background represent the individual realizations of the bestfit model. For the data (shown in black dashed lines), we show both optimistic and pessimistic limits, although we used only the optimistic case for the parameter space exploration. We checked that the bestfit evolution model can nicely match the observed scatter in the optical depth distribution as well as the mean transmission flux in the majority of the cases with mild deviation at $z=5.6$ and $5.9$. The deviation is mainly due to the dominance of $\tau_{\mathrm{eff}}$ values at the distribution of the edges in the data for those redshifts. Furthermore, the width of the distribution increases as we go towards higher redshifts, suggesting increasing fluctuations in the opacity fields.
@@ 5.2 Joint parameter posteriors
Next, we quantify the posterior distribution of the parameters extracted from three different temperature models allowed by the current constraints on the IGM temperatures, as shown earlier in Figure 5. We perform this analysis for all the redshifts of interests and show the combined posteriors in Figure 11 while in Figure 12, we show the 2D posterior contours (at 68% and 95% confidence level) for the individual temperature models along with the combined contour of all the three samples again at the fiducial redshift, $z=5.4$. It is apparent that the combined contour nicely represents the whole posterior space allowed by the temperature constraints. Now, moving the discussion to Figure 11 for different redshifts, it can be seen that the data are consistent even with almost uniform UVB models towards low redshifts ($z\leq 5.3$), demanding only a small amount of UVB fluctuations. This interpretation is driven by the fact that, in our definition, post-reionization UVB fluctuations are solely quantified by the mean free path parameter $\lambda_{0}$. The large parameter values correspond to a relatively uniform UVB field and vice versa. As at those redshifts, all the posteriors hit the upper end of the prior values on $\lambda_{0}$, we can conclude that almost uniform UVB models are also allowed by the data. An interesting feature can be noticed at $z=5.3$ where a bimodality in mean free path distribution arises disfavouring some of the moderate UVB fluctuation models ($\lambda_{0}\sim 20-30h^{-1}\mathrm{cMpc}$). This happens because those models are not able to provide enough fluctuations to match the data distribution. For a large mean free path, the opacity fluctuations are driven by the underlying density field and temperature fluctuations. As the mean free path decreases, it initially neutralizes the opacity fluctuations as demonstrated in Figure 7, which corresponds to the disfavored region in the posterior. This happens because high-density regions are ionized earlier and have relatively lower temperatures at later times due to the availability of more time for cooling (D’Aloisio et al., 2015). This enhances the opacity fluctuations correlating with density fluctuations. However, the correlation between density and UVB field is opposite, providing high photoionization rate at high density regions and vice versa, which further drives the fluctuations in the other direction (see also; Davies et al., 2018a). Then, with a reasonably short mean free path, the opacity fluctuations are dominated by the UVB field itself. Hence, the bimodal feature portrays the interplay among these competing fluctuations, generating the opacity field.
@@
As we move towards higher redshifts (up to $z=5.8$, where reionization ends in our model), the required UVB fluctuations also increase, allowing for shorter mean free paths. Before reionization ends ($z\geq 5.9$), the ionization fluctuations (driven by leftover neutral islands) are strong enough to produce the required large-scale UVB fluctuations matching the observed opacity fluctuations, hence requiring little or almost no fluctuations driven by post-reionization $\lambda_{0}$. This further keeps the mean free path parameter unconstrained. This explanation can be clearer from Figure 13, where we show two snapshots of UVB fluctuations. Among these two snapshots, one is at redshift, $z=5.8$ (after reionization completion) with a very short mean free path ($\lambda_{0}=4\penalty 10000\ h^{-1}\mathrm{cMpc}$) and the other one is at redshift, $z=5.9$ (before reionization completion) with a relatively high mean free path parameter ($\lambda_{0}=40\penalty 10000\ h^{-1}\mathrm{cMpc}$), as suggested by the posteriors. It is apparent that the presence of neutral regions can produce strong fluctuations in the UVB field even if $\lambda_{0}$ is larger (see also, Choudhury et al., 2021). This supports our above argument for an unconstrained mean free path at higher redshifts where reionization is still not ended. Hence, despite the fact that the $\lambda_{0}$ posteriors look very similar to the cases at $z\leq 5.3$, the UVB fluctuations are not uniform here (at z=5.96.1z=5.9-6.1), rather driven by the leftover neutral islands. This analysis provides us with the confidence that our simplistic model can be utilized to pursue parameter space statistics study and interpret the relevant physics from the observational data.
@@
In Figure 14, we show the marginalized constraints on the mean photoionization rate ($\langle\Gamma_{12}^{\mathrm{HI}}\rangle$, at the left panel) and effective mean free path ($\lambda_{0}$, at the right panel) at 95% confidence level along with the points from recent literature. Our estimates are denoted with red circles (the solid ones indicate the redshifts where reionization is completed, and hollow ones indicate the redshifts with incomplete reionization). We found that our photoionization rate estimates are consistent with the findings of Davies et al. (2024b, shown in blue stars with displaced redshift for better visualization). This is not very surprising, as the UVB fluctuation generation model is the same for both cases. However, our points are slightly higher as we have a realistic reionization history allowing sufficient time for cooling in the early ionized high-density regions. On the other hand, Davies et al. (2024b) assumed a sudden reionization for all the cells in the simulation box, which results in larger temperature and hence, a slightly shorter photoionization rate to match the mean transmission flux. Additionally, our estimates are also in agreement with the corresponding constraints of Gaikwad et al. (2023, shown in green triangle again shifted in redshift).
@@
In the case of the mean free path, we are able to derive the constraints from posteriors at redshift range z=5.45.7z=5.4-5.7. For the redshifts providing constraints from only one side, we quote the value where the posterior value falls down to $1/e^{2}$ from the maximum (Davies et al., 2024b). We do not get any constraints at $z=6.0$. Our constraints prefer a shorter mean free path distribution than Davies et al. (2024b) values, although these are consistent within the error bars, particularly given that the lower end of the Davies et al. (2024b) uncertainties corresponds to an approximate correction to account for their lack of model self-consistency. The main reason for this difference is again traced back to the assumed sudden reionization and the lack of post-reionization large-scale temperature fluctuations in their model. It is expected that the large-scale temperature field is anticorrelated with the density field due to the variations in reionization timing at different regions. As a result of that, a stronger UVB fluctuation is required to overcome the correlated fluctuations of thermal and density field, which in turn demands a lower $\lambda_{0}$ value. Intriguingly, this more physical treatment with a gradual and patchy thermal evolution puts our model constraints within a similar ballpark of recent high-redshift mean free path measurements (Becker et al., 2021; Zhu et al., 2023). At this point, it is important to highlight that the definition of our mean free path is similar to Davies et al. (2024b), which is solely a subgrid parameter quantifying clumping of the unresolved gas. Hence, our inferred mean free path only characterizes the excess UVB fluctuations over the uniform case without including density field fluctuations. As a consequence, we do not get any upper bound on $\lambda_{0}$ for $z\leq 5.3$ as an almost uniform UVB field is consistent with fluctuations required by the data. This is unlike the approach adopted by Gaikwad et al. (2023), where the mean free path is computed along with their density field after including fluctuating UVB contribution. That further enables them to put constraints on the mean free path even without requiring explicit UVB fluctuations. At redshift $z\geq 5.9$, reionization is incomplete in our model, which further allows us to provide only lower limits on the mean free path, as demonstrated earlier in the subsection 5.2. The constraints used in the plots are tabulated in Table 1.
@@
In continuation with the above discussion, it is worthwhile to mention that most of the semi-numerical models give a sharp transition to the reionization end, while detailed numerical simulations (Kannan et al., 2022; Asthana et al., 2025) prefer relatively smooth transitions along with slight inflection in neutral fraction evolution. A gradual transition will translate to a smoother evolution in mean free path, as hinted from our findings. A smooth transition towards reionization end can be achieved within a semi-numerical setup with more realistic modelling of photon sinks (e.g. Davies & Furlanetto 2022; Qin et al. 2025). This can provide a direction for potential improvement for our model in the future.
@@
Table 1: Parameter constraints obtained from the MCMC-based analysis after combining different temperature models. For each redshift, we show the mean value of combined samples along with 95% confidence limits on different parameters where constraints are achieved from both sides. We quote the value where the posterior value falls down to $1/e^{2}$ from the maximum (Davies et al., 2024b) for one sided limit.